# Preference Optimization by Estimating the Ratio of the Data Distribution

### v1

@@ 3.4 Compatibility with other DPO extensions
where $f:\mathbb{R}^{+}\rightarrow\mathbb{R}$ denotes a convex function satisfying $f(1)=0$. Substituting $R_{\theta}^{f\text{-DPO}}$ and the $h$ function corresponding to logistic regression into BPO; $\mathcal{L}^{\text{LR}}_{\text{BPO}}(R_{\theta}^{f\text{-DPO}};p_{\text{data}})$ recovers the original objective of $f$-DPO wang2024beyond (See Section A.6 for details). BPO enables constructing new objectives by varying the choice of $h$ based on the model ratio $R_{\theta}^{f\text{-DPO}}$, which results in $\mathcal{L}^{h}_{\text{BPO}}(R_{\theta}^{f\text{-DPO}};p_{\text{data}})$. Similarly, existing logistic regression-based DPO variants can also be generalized by BPO.
Table 3: Comparison of dialogue generation performance on Anthropic-HH, using the same SFT model based on Pythia-2.8B. Each value is colored red if it is better than DPO, and blue if it performs worse. The best result for each metric is shown in bold, and the second and third best are underlined.
@@
4 Experiments
This section presents the empirical performance of the proposed BPO compared to prior preference optimization methods. Section 4.1 compares instances of BPO with other probabilistic loss extensions. Section 4.2 compares BPO to the state-of-the-art DPO loss variants on popular LLM benchmarks.
4.1 Comparison with probabilistic DPO loss extensions
Task & experimental setup The experiments are conducted for single-turn dialogue generation using the Anthropic helpful and harmless (HH) dataset bai2022training , and summarization using the Reddit TL;DR dataset volske2017tl . While the experimental setups vary across the different baseline papers, we faithfully follow the setup provided in the original DPO paper rafailov2023direct for a fair comparison. For dialogue generation, we use Pythia-2.8B biderman2023pythia as the pre-trained LLM and perform SFT on the preferred subset of the HH dataset. For summarization, we use a publicly available SFT model tldr-sft based on GPT-J gpt-j . All comparisons are conducted on the same SFT model with identical training hyperparameters ($\beta$, learning rate, batch size). See supplementary materials for details.
Baselines We compare with representative baselines, including instances of $f$-DPO wang2024beyond and $f$-PO xu2025fpo . For $f$-DPO, we consider forward KL (FKL), Jensen-Shannon (JS), $\alpha$-divergence with $\alpha=0.1,0.3,0.5,$ and $0.7$, selecting $\alpha=0.1$ based on the best win rate, and additionally include $\chi^{2}$-divergence huang2025correcting . For $f$-PO, we consider Jeffrey, JS, reverse KL (RKL) ji2024towards , and $\alpha$-divergence with $\alpha=0.1$, as suggested in the original paper. To ensure a fair computational comparison, we adopt the pairwise approximation proposed in $f$-PO.
Evaluation metrics Both fidelity and diversity of the generated samples are important from a probability matching perspective. We evaluated the models using test datasets. To measure fidelity, we use the win rate against both the preferred responses and the SFT model responses, as judged by GPT-4. To assess diversity, we use predictive entropy zengtoken ; wang2024beyond , self-BLEU zhu2018texygen , and distinct-1 li2016diversity .
@@ 4.1.1 Dialogue generation task
Figure 5: Ablation studies on the effect of $\lambda$ in SBA, using Anthropic-HH with Pythia-2.8B.
Comparison to baselines Figure 1 shows the trade-off between win rate and entropy for each instance, and Table 3 summarizes the results across all metrics. $f$-DPO recovers DPO when RKL is used as the regularizer. As the $\alpha$-divergence shifts from RKL to FKL, the win rate decreases while diversity improves, reflecting the mode-covering nature of FKL. $\chi^{2}$ improves win rate, as observed in $\chi^{2}$-PO huang2006correcting . $f$-PO recovers DPO with FKL. Other divergences, such as Jeffrey, JS, $\alpha$-divergence, and RKL, have weaker mode coverage than FKL, resulting in worse diversity across all metrics.
@@
Orthogonal utilization with other methods As discussed in Section 3.4, existing loss functions that can be expressed as logistic regression can be orthogonally extended using BPO. Table 4 reports the performance of BPO applied to two instances of $f$-DPO (FKL and JS), showing improvements in 9 out of 10 metrics. As shown in Section B.2, $f$-PO with pairwise datasets can also be viewed as logistic regression for some $f$, and applying SBA to $f$-PO (JS) results in further performance gains. Figure 4 summarizes the performance improvements achieved by applying SBA to each model ratio.
Ablation studies on $\lambda$ in SBA Figure 5 shows ablation studies on the effect of $\lambda$ in the proposed SBA, an instance of the BPO framework. When $\lambda=-0.5$, the gradient behavior closely resembles that of DPO, as shown in Figure 3(b), with both win rate and entropy showing similar trends in Figures 5(a) and 5(b). As $\lambda$ increases, both metrics improve up to a certain point, in contrast to the instances of $f$-PO and $f$-DPO that exhibit a clear trade-off between win rate and entropy. Larger values of $\lambda$ lead the model to focus more on confident samples (where $R_{\theta}$ deviates further from 1), resulting in a wider spread in reward margin ($-\log R_{\theta}$) statistics, as shown in Figure 5(c). This shift in the statistics of reward margins appears to be an underlying factor contributing to the performance improvements.
@@ 4.1.2 Summarization task
For the TL;DR summarization task, we compare the best win rate instances of each framework reported in Table 3, where $f$-DPO uses $\chi^{2}$, $f$-PO uses JS, and BPO uses SBA. Figure 6 presents the win rate against preferred responses across different sampling temperatures, showing a trend consistent with prior studies that lower temperatures lead to better performance. BPO consistently outperforms the other instances at all temperatures. Table 5 shows additional metrics at temperature 0.25, where BPO shows the best performance across all metrics.
Table 5: Performance on TL;DL summarization based on GPT-J with 0.25 temperature.
@@
4.2 Comparison with general DPO loss variants
Experimental setup We conduct experiments using Mistral-7B jiang2023mistral7b and Llama-3-8B-Instruct grattafiori2024llama backbone models, based on the UltraFeedback dataset cui2024ultrafeedback . For Mistral-7B, we use publicly available SFT models from Zephyr tunstall2023zephyr ; zephyr-7b-sft-full . We extend DPO with our BPO loss based on the Zephyr setup and extend SimPO with our BPO loss following the SimPO meng2024simpo setup. For Llama-3-8B-Instruct, we use the same SFT models Meta-Llama-3-8B-Instruct as in the SimPO setup, along with the adapted version of the UltraFeedback dataset llama3-ultrafeedback-armorm for Llama-3-8B-Instruct. See supplementary materials for details.
Evaluation benchmark We evaluate the response capability across a wide range of queries using the most popular open-ended instruction-following benchmarks. AlpacaEval2 alpaca_eval measures the win rate on 805 examples, using GPT-4 Turbo as both the judge and the opponent. The evaluation covers both raw and length-controlled responses. Arena-Hard li2024crowdsourced measures the win rate on 500 queries, using GPT-4 Turbo as the judge and GPT-4-0314 as the opponent. All evaluations follow the library configurations and decoding parameters used in SimPO to ensure fair comparisons.
Table 6: Performance on AlpacaEval2 and Arena-Hard with Llama-3-8B-Instruct. LC indicates the length-controlled win rate, and WR denotes the raw win rate.
Method AlpacaEval 2.0 Arena-Hard
LC (%) WR (%) WR (%)
SFT 26.0 25.3 22.3
SimPO-v1 meng2024simpo 44.7 40.5 33.8
$f$-PO xu2025fpo 45.3 41.0 33.5
RRHF yuan2023rrhf 37.9 31.6 28.8
SLiC-HF zhao2023slic 33.9 32.5 29.3
DPO rafailov2023direct 48.2 47.5 35.2
IPO azar2024general 46.8 42.4 36.6
CPO pmlr-v235-xu24t 34.1 36.4 30.9
KTO pmlr-v235-ethayarajh24a 34.1 32.1 27.3
ORPO hong2024orpo 38.1 33.8 28.2
R-DPO park2024disentangling 48.0 45.8 35.1
SimPO-v2 meng2024simpo 53.7 47.5 36.5
BPO 55.9 51.5 38.0
Results Table 6 presents the main results using Llama-3-8B-Instruct. Baseline results are taken from SimPO meng2024simpo and $f$-PO xu2025fpo . Note that both SimPO-v1 and $f$-PO are trained on an older version of the datasets llama3-ultrafeedback . The $f$-PO model trained on top of SimPO-v1 shows less than 1% win rate improvement on both LC and WR in AlpacaEval2, and even performs worse on Arena-Hard. All other baselines and BPO are trained on the same dataset as SimPO-v2. Most DPO variants show even worse performance than standard DPO. BPO can also be applied on top of SimPO-v2 as discussed in Section 3.4, and BPO achieves a 2.2% gain in LC, a 4.0% gain in WR, and a 1.5% improvement on Arena-Hard compared to SimPO-v2. While $f$-PO built on SimPO-v1 achieved only marginal improvements, BPO provides substantial gains. To the best of our knowledge, BPO achieves state-of-the-art LC performance among preference-optimized models based on Llama-3-8B-Instruct. Figure 7 shows results with Mistral-7B. Specifically, BPO improves LC from 15.1% to 19.1% when built on DPO, and from 21.5% to 23.7% when built on SimPO. These results demonstrate that BPO generalizes well to external benchmarks beyond the training domain with the 7–8B LLM scale, showing consistent improvements across different backbones and model ratio formulations, including DPO and SimPO.
@@ 5 Related work
DPO loss generalization in distribution matching perspective $f$-DPO wang2024beyond generalizes the reverse KL regularization in $\mathcal{L}_{\text{RLHF}}$ (see Eq. 2) to the broader class of $f$-divergences csiszar2004information ; liese2006divergences , and deriving direct optimization objectives for a subset of $f$-divergences that satisfy specific conditions. $\chi^{2}$-PO huang2025correcting extends the direct optimization objective for $\chi^{2}$ divergence regularization within the $f$-DPO framework. However, modifying the regularization alone is potentially insufficient to determine the overall behavior of the loss function and naturally leads to a different optimal policy, as we discussed in Section B.1. EXO ji2024towards reformulates the entire DPO objective as a distribution matching problem, showing that $\mathcal{L}_{\text{DPO}}$ in Eq. 4 reduces to the forward KL divergence between the policy model $\pi_{\theta}({\mathbf{y}}|{\mathbf{x}})$ and the optimal policy $\pi_{\theta^{*}}({\mathbf{y}}|{\mathbf{x}})$ in Eq. 3. EXO further proposes a reverse KL loss, and $f$-PO xu2025fpo subsequently extends this formulation to $f$-divergences. In contrast to $f$-DPO, which yields a different optimal policy, $f$-PO is analogous to extensions of traditional generative modeling nowozin2016f ; song2021train that directly match a policy to a target optimal policy. However, the $f$-PO loss requires a learned reward model $r_{\phi^{*}}$ and Monte Carlo estimation of partition functions to compute the probability of the optimal policy $\pi_{\theta^{*}}$. Simplifying $f$-PO for a paired preference dataset results in a different notion of target optimality as we discussed in Section B.2. As summarized in Table 1, the proposed BPO is distinguished from existing probabilistic loss extensions by maintaining both optimality and simplicity, while allowing flexible choices of optimization objectives.
@@ Appendix A Proofs and derivations
$p_{\text{data}}({\mathbf{y}}_{w}\succ{\mathbf{y}}_{l}\mid{\mathbf{x}}),\pi_{\text{ref}}({\mathbf{y}}\mid{\mathbf{x}}),\pi_{\theta^{*}}({\mathbf{y}}\mid{\mathbf{x}})>0$ for all ${\mathbf{y}}_{w},{\mathbf{y}}_{l},{\mathbf{y}},{\mathbf{x}}$
@@
According to the definition of Bregman divergence [7], the function $h$ is required to be strictly convex and continuously differentiable. Since we optimize over this divergence, an additional condition on its second derivative is needed. This condition is required in previous works [25] that use Bregman divergence as a loss. The well-known Bregman divergences listed in Table 2 satisfy the condition.

### v2

@@ 3.4 Compatibility with other DPO extensions
where $f:\mathbb{R}^{+}\rightarrow\mathbb{R}$ denotes a convex function satisfying $f(1)=0$. Substituting $R_{\theta}^{f\text{-DPO}}$ and the $h$ function corresponding to logistic regression into BPO; $\mathcal{L}^{\text{LR}}_{\text{BPO}}(R_{\theta}^{f\text{-DPO}};p_{\text{data}})$ recovers the original objective of $f$-DPO wang2024beyond (See Section˜A.6 for details). BPO enables constructing new objectives by varying the choice of $h$ based on the model ratio $R_{\theta}^{f\text{-DPO}}$, which results in $\mathcal{L}^{h}_{\text{BPO}}(R_{\theta}^{f\text{-DPO}};p_{\text{data}})$. Similarly, existing logistic regression-based DPO variants can also be generalized by BPO.
4 Experiments
This section presents the empirical performance of the proposed BPO compared to prior preference optimization methods. Section˜4.1 compares instances of BPO with other probabilistic loss extensions. Section˜4.2 compares BPO to the state-of-the-art DPO loss variants on popular LLM benchmarks.
4.1 Comparison with probabilistic DPO loss extensions
Task & experimental setup: The experiments are conducted for single-turn dialogue generation using the Anthropic helpful and harmless (HH) dataset bai2022training , and summarization using the Reddit TL;DR dataset volske2017tl . While the experimental setups vary across the different baseline papers, we faithfully follow the setup provided in the original DPO paper rafailov2023direct for a fair comparison. For dialogue generation, we use Pythia-2.8B biderman2023pythia as the pre-trained LLM and perform SFT on the preferred subset of the HH dataset. For summarization, we use a publicly available SFT model tldr-sft based on GPT-J gpt-j . All comparisons are conducted on the same SFT model with identical training hyperparameters ($\beta$, learning rate, batch size). See Appendix˜C for details.
Table 3: Comparison of dialogue generation performance on Anthropic-$H$, using the same SFT model based on Pythia-2.8B. Each value is colored red if it is better than DPO, and blue if it performs worse. The best result for each metric is shown in bold, and the second and third best are underlined.
@@ 4.1 Comparison with probabilistic DPO loss extensions
Baselines: We compare with representative baselines, including instances of $f$-DPO wang2024beyond and $f$-PO xu2025fpo . For $f$-DPO, we consider forward KL (FKL), Jensen-Shannon (JS), $\alpha$-divergence with $\alpha=0.1,0.3,0.5,$ and $0.7$, selecting $\alpha=0.1$ based on the best win rate, and additionally include $\chi^{2}$-divergence huang2025correcting . For $f$-PO, we consider Jeffrey, JS, reverse KL (RKL) ji2024towards , and $\alpha$-divergence with $\alpha=0.1$, as suggested in the original paper. To ensure a fair computational comparison, we adopt the pairwise approximation proposed in $f$-PO.
Evaluation metrics: Both fidelity and diversity of the generated samples are important from a probability matching perspective. We evaluated the models using test datasets. To measure fidelity, we use the win rate against both the preferred responses and the SFT model responses, as judged by GPT-4. To assess diversity, we use predictive entropy zengtoken ; wang2024beyond , self-BLEU zhu2018texygen , and distinct-1 li2016diversity .
@@ 4.1.1 Dialogue generation task
Figure 5: Ablation studies on the effect of $\lambda$ in SBA, using Anthropic-$H$ with Pythia-2.8B.
Comparison to baselines: Figure˜1 shows the trade-off between win rate and entropy for each instance, and Table˜3 summarizes the results across all metrics. $f$-DPO recovers DPO when RKL is used as the regularizer. As the $\alpha$-divergence shifts from RKL to FKL, the win rate decreases while diversity improves, reflecting the mode-covering nature of FKL. $\chi^{2}$ improves win rate, as observed in $\chi^{2}$-PO huang2006correcting . $f$-PO recovers DPO with FKL. Other divergences, such as Jeffrey, JS, $\alpha$-divergence, and RKL, have weaker mode coverage than FKL, resulting in worse diversity across all metrics.
@@
Orthogonal utilization with other methods: As discussed in Section˜3.4, existing loss functions that can be expressed as logistic regression can be orthogonally extended using BPO. Table˜4 reports the performance of BPO applied to two instances of $f$-DPO (FKL and JS), showing improvements in 9 out of 10 metrics. As shown in Section˜B.2, $f$-PO with pairwise datasets can also be viewed as logistic regression for some $f$, and applying SBA to $f$-PO (JS) results in further performance gains. Figure˜4 summarizes the performance improvements achieved by applying SBA to each model ratio.
Ablation studies on $\lambda$ in SBA: Figure˜5 shows ablation studies on the effect of $\lambda$ in the proposed SBA, an instance of the BPO framework. When $\lambda=-0.5$, the gradient behavior closely resembles that of DPO, as shown in Figure˜3(b), with both win rate and entropy showing similar trends in Figures˜5(a) and 5(b). As $\lambda$ increases, both metrics improve up to a certain point, in contrast to the instances of $f$-PO and $f$-DPO that exhibit a clear trade-off between win rate and entropy. Larger values of $\lambda$ lead the model to focus more on confident samples (where $R_{\theta}$ deviates further from 1), resulting in a wider spread in reward margin ($\log R_{\theta}$) statistics, as shown in Figure˜5(c). This shift in the statistics of reward margins appears to be an underlying factor contributing to the performance improvements.
@@ 4.1.2 Summarization task
For the TL;DR summarization task, we compare the best win rate instances of each framework reported in Table˜3, where $f$-DPO uses $\chi^{2}$, $f$-PO uses JS, and BPO uses SBA. Figure˜6 presents the win rate against preferred responses across different sampling temperatures, showing a trend consistent with prior studies that lower temperatures lead to better performance. BPO consistently outperforms the other instances at all temperatures. Table˜5 shows additional metrics at temperature 0.25, where BPO shows the best performance across all metrics.
4.2 Comparison with general DPO loss variants
Experimental setup: We conduct experiments using Mistral-7B-Base jiang2023mistral7b , Llama-3-8B-Base, and Llama-3-8B-Instruct grattafiori2024llama backbone models, based on the UltraFeedback dataset cui2024ultrafeedback . For Mistral-7B-Base, we use publicly available SFT models from Zephyr tunstall2023zephyr ; zephyr-7b-sft-full . We extend DPO with our BPO loss based on the Zephyr setup and extend SimPO with our BPO loss following the SimPO meng2024simpo setup. For Llama-3-8B-Base and Llama-3-8B-Instruct, we use the same SFT models Meta-Llama-3-8B-Instruct as in the SimPO setup, along with the adapted version of the UltraFeedback dataset llama3-ultrafeedback-armorm . See Appendix˜C for details.
Evaluation benchmark: We evaluate the response capability across a wide range of queries using the most popular open-ended instruction-following benchmarks. AlpacaEval2 alpaca_eval measures the win rate on 805 examples, using GPT-4 Turbo as both the judge and the opponent. The evaluation covers both raw and length-controlled responses. Arena-Hard li2024crowdsourced measures the win rate on 500 queries, using GPT-4 Turbo as the judge and GPT-4-0314 as the opponent. All evaluations follow the library configurations and decoding parameters used in SimPO to ensure fair comparisons.
Table 5: Performance on TL;DR summarization based on GPT-J with 0.25 temperature.
@@ 4.2 Comparison with general DPO loss variants
Table 6: Performance on AlpacaEval2 and Arena-Hard with various backbone models. LC (%) indicates the length-controlled win rate, and WR (%) denotes the raw win rate. The baseline results are from SimPO meng2024simpo .
Mistral-7B-Base Llama-3-8B-Base Llama-3-8B-Instruct
Method AlpacaEval Arena-Hard AlpacaEval Arena-Hard AlpacaEval Arena-Hard
LC WR WR LC WR WR LC WR WR
SFT 8.4 6.2 1.3 6.2 4.6 3.3 26.0 25.3 22.3
RRHF yuan2023rrhf 11.6 10.2 5.8 12.1 10.1 6.3 37.9 31.6 28.8
SLiC-HF zhao2023slic 10.9 8.9 7.3 12.3 13.7 6.0 33.9 32.5 29.3
DPO rafailov2023direct 15.1 12.5 10.4 18.2 15.5 15.9 48.2 47.5 35.2
IPO azar2024general 11.8 9.4 7.5 14.4 14.2 17.8 46.8 42.4 36.6
CPO pmlr-v235-xu24t 9.8 8.9 6.9 10.8 8.1 5.8 34.1 36.4 30.9
KTO pmlr-v235-ethayarajh24a 13.1 9.1 5.6 14.2 12.4 12.5 34.1 32.1 27.3
ORPO hong2024orpo 14.7 12.2 7.0 12.2 10.6 10.8 38.1 33.8 28.2
R-DPO park2024disentangling 17.4 12.8 8.0 17.6 14.4 17.2 48.0 45.8 35.1
SimPO meng2024simpo 21.5 20.8 16.6 22.0 20.3 23.4 53.7 47.5 36.5
BPO 23.7 20.9 16.9 22.5 18.7 31.7 55.9 51.5 38.0
Results: Table˜6 presents the main results obtained with various backbone models. Baseline results are taken from SimPO meng2024simpo . Most DPO variants exhibit even worse performance than the standard DPO. BPO can also be applied on top of SimPO, as discussed in Section˜3.4, and it achieves consistent performance gains over SimPO, except in one out of nine cases. To the best of our knowledge, BPO achieves state-of-the-art LC performance among preference-optimized models based on Llama-3-8B-Instruct. Figure˜7 shows results with Mistral-7B-Base. Specifically, BPO improves LC from 15.1% to 19.1% when built on DPO, and from 21.5% to 23.7% when built on SimPO. These results demonstrate that BPO generalizes well to external benchmarks beyond the training domain with the 7–8B LLM scale, showing consistent improvements across different backbones and model ratio formulations, including DPO and SimPO.
@@ 5 Related work
DPO loss generalization in distribution matching perspective: $f$-DPO wang2024beyond generalizes the reverse KL regularization in $\mathcal{L}_{\text{RLHF}}$ (see Eq.˜2) to the broader class of $f$-divergences csiszar2004information ; liese2006divergences , and deriving direct optimization objectives for a subset of $f$-divergences that satisfy specific conditions. $\chi^{2}$-PO huang2025correcting extends the direct optimization objective for $\chi^{2}$ divergence regularization within the $f$-DPO framework. MPO alfano2024meta generalizes the regularization under the Bregman divergence, particularly for meta-learning. However, modifying the regularization alone is potentially insufficient to determine the overall behavior of the loss function and naturally leads to a different optimal policy, as we discussed in Section˜B.1. EXO ji2024towards reformulates the entire DPO objective as a distribution matching problem, showing that $\mathcal{L}_{\text{DPO}}$ in Eq.˜4 reduces to the forward KL divergence between the policy model $\pi_{\theta}({\mathbf{y}}|{\mathbf{x}})$ and the optimal policy $\pi_{\theta^{*}}({\mathbf{y}}|{\mathbf{x}})$ in Eq.˜3. EXO further proposes a reverse KL loss, and $f$-PO xu2025fpo subsequently extends this formulation to $f$-divergences. In contrast to $f$-DPO, which yields a different optimal policy, $f$-PO is analogous to extensions of traditional generative modeling nowozin2016f ; song2021train that directly match a policy to a target optimal policy. However, the $f$-PO loss requires a learned reward model $r_{\phi^{*}}$ and Monte Carlo estimation of partition functions to compute the probability of the optimal policy $\pi_{\theta^{*}}$. Simplifying $f$-PO for a paired preference dataset results in a different notion of target optimality as we discussed in Section˜B.2. As summarized in Table˜1, the proposed BPO is distinguished from existing probabilistic loss extensions by maintaining both optimality and simplicity, while allowing flexible choices of optimization objectives.
Generative modeling and likelihood ratio estimation: The likelihood ratio has played a central role in research on generative modeling. For continuous random variables, noise contrastive estimation (NCE) gutmann2012noise estimates the ratio between a known noise distribution and the target distribution to approximate the target density. This ratio identifies the target density and also serves as a learning signal for neural samplers, inspiring subsequent work on GANs goodfellow2014generative . The noise distribution in NCE has been replaced by the model distributions of GANs azadi2018discriminator ; che2020your , VAEs makhzani2015adversarial ; aneja2021contrastive , and diffusion models pmlr-v202-kim23i ; na2024diffusion , and has recently been employed as a refinement technique in generative modeling. For discrete variables, concrete score matching (CSM) meng2022concrete estimates the ratio of probability masses between different states of a probability distribution. CSM has served as a foundation for the development of discrete diffusion modeling lou2024discrete ; zhang2025target , and this paper adopts the concept of CSM to autoregressive language models. Another concurrent work has utilized CSM for knowledge distillation in autoregressive language models kim2025distillation .
@@ 6 Conclusion
Acknowledgment
This work was supported by the IITP (Institute of Information & Communications Technology Planning & Evaluation)-ITRC (Information Technology Research Center) grant funded by the Korea government (Ministry of Science and ICT) (IITP-2025-RS-2024-00437268).
@@ Appendix A Proofs and derivations
$p_{\text{data}}({\mathbf{y}}_{w}\succ{\mathbf{y}}_{l}\mid{\mathbf{x}}),\pi_{\text{ref}}({\mathbf{y}}\mid{\mathbf{x}}),\pi_{\theta^{*}}({\mathbf{y}}\mid{\mathbf{x}})>0$ for all ${\mathbf{y}}_{w},{\mathbf{y}}_{l},{\mathbf{y}},{\mathbf{x}}$ .
@@
According to the definition of Bregman divergence [10], the function $h$ is required to be strictly convex and continuously differentiable. Since we optimize over this divergence, an additional condition on its second derivative is needed. This condition is required in previous works [30] that use Bregman divergence as a loss. The well-known Bregman divergences listed in Table˜2 satisfy the condition.